OpenStudy (anonymous):

2. solve the triangle below by finding all missing sides and or angles part1: find the length of side AC. round your answer to nearest hundreths part2: find the measure of angle C part3: find the measure of angle A using any method

OpenStudy (anonymous):



OpenStudy (anonymous):

Use the law of cosines

OpenStudy (anonymous):

AC^2=AB^2+BC^2-2*BC*AB*cos(B)

OpenStudy (anonymous):

You can then use the law of sines to solve the rest of it

OpenStudy (dumbcow):

yes use law of cosines to find AC then using law of sines sin(C)/10 = sin(110)/AC

OpenStudy (anonymous):

how do i find the AC BC AB in order to use the formula tho?

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

AB = 10 BC = 8 angle B = 110

OpenStudy (anonymous):

well duh tdiaz100, lol okay let me solve it

OpenStudy (dumbcow):

haha ok:)

OpenStudy (anonymous):

323.84

OpenStudy (anonymous):

????

OpenStudy (dumbcow):

for AC, no

OpenStudy (anonymous):

ok i know question 2 and 3 is 35 degrees

OpenStudy (anonymous):

Take the square root

OpenStudy (dumbcow):

it has to be less than 18

OpenStudy (anonymous):

17.99

OpenStudy (anonymous):

i round to the nearest hundreths would it be 20?

OpenStudy (dumbcow):

something went wrong =100+64-2(80)cos(110)

OpenStudy (anonymous):

were the 80 came from?

OpenStudy (dumbcow):

10*8

OpenStudy (anonymous):

same thing 17.99

OpenStudy (anonymous):

14.79

OpenStudy (dumbcow):

no that would mean the angle is like 179 degrees

OpenStudy (anonymous):

is your calculator in degree mode?

OpenStudy (dumbcow):

haha yeah thats what it is

OpenStudy (anonymous):

yes it was

OpenStudy (anonymous):

If i put mine in radian mode, i get 17.9956. in degree mode, i get 14.7893

OpenStudy (anonymous):

the nearest hundeths is 14.89 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

nt 79

OpenStudy (anonymous):

14.79*

OpenStudy (anonymous):

yes 14.79 srry